f of g(x) and g of f(x)where f(x)=x2 the x is supposed to be squared and g(x)= x-3 next part of same problem is the domain of f of g and the domain of g of f
Answer by tinbar(133)

f(g(x)) = f(x-3) = (x-3)^2 = x^2-6x+9
g(f(x)) = g(x^2) = (x^2)-3
domain for f(g(x)): First we need the domain for g(x) which is all Real numbers since we can subtract 3 from any x in R(set of Real Numbers). Next, out of all these numbers, we can square any of them, therefore the domain of f(g(x)) is the entire set of Real numbers
domain for g(f(x)): Similar argument as the first one, resulting in the same domain